OpenStudy (anonymous):

Find the distance between points P(4, 6) and Q(8, 9) to the nearest tenth.

OpenStudy (anonymous):

Distance between two points P(4, 6) and Q(8, 9) can be found from\[\sqrt{(8-4)^2+(9-6)^2}=\sqrt{25}=5cm\]